A 0.300-kg piece of iron initially at 315 ºC is dropped into 0.400 L of water initially at 530ºR. Assuming that all heat transfer occurs between the iron and the water, calculate the final temperature.


1
Expert's answer
2021-01-20T03:14:00-0500

The heat change of water is equal to the heat change of iron.

So,

"0.3*1000*0.45(273+315-T)=400*4.18*(T-(530+273))\newline
135(588-T)=180(T-803)\newline 144540-79380=45T\newline 65160=45T\newline T=1448K \newline The\ final\ temperature\ is\ 1448k" "In\ degrees\ =1175degrees"